OpenStudy (anonymous):

what is the square root of 75y^4w^3

OpenStudy (vijay):

\[5y ^{2}w \sqrt{3w}\]

OpenStudy (anonymous):

can you please explain it?

OpenStudy (vijay):

75 = 3 * 5^2 y^4 = (y^2)^2 w^3 = w * w^2

OpenStudy (vijay):

then when u put squareroot, all power 2 will be ignore

OpenStudy (vijay):

put all other balance into the sqrt